Shaypop:

Draw a square ABCD with diagonals AD and BC. Label the point intersection E then use the following information to find the measure. AB= 6x-3 BD= 4x+5 BC=16 AB BD AC AD CE AE m

Shaypop:

Can somebody help me

Shaypop:

Can somebody draw that for me in the comments

Angle:

AB = BC because a square has equal side lengths this means 6x - 3 = 16 solve for x

Angle:

once you get the value for x, you can plug that into AB = 6x - 3 and BD = 4x + 5 to get those answers

Angle:

BD = AC

Angle:

AD = AB

Angle:

CE = BD divided by 2 AE = BD divided by 2

Angle:

angle BAC = 90 divided by 2 angle BDC = 90 divided by 2

Angle:

that is all of the information you need to be able to get your answers.

Shaypop:

How is it 90

Angle:

the corners of a square are always 90

Shaypop:

The value of x is 3.17

Angle:

yes

Shaypop:

You doing it all wrong

Eiwoh2:

I’m sure a literal math teacher knows what they’re doing lol

Angle:

cool, so if you keep being a brat about it - you won't get any more help :)

Eiwoh2:

You’re the one needing help, correct?

Shaypop:

Yes I need help

Eiwoh2:

Then accept it, Angle is guiding you her way to helping you understand your question. If you say she isn’t doing it right, what’s the point of asking for help?

Shaypop:

Im sorry

Shaypop:

Angle so put 3.17 to what to plug in in

Angle:

\(\color{#0cbb34}{\text{Originally Posted by}}\) Angle once you get the value for x, you can plug that into AB = 6x - 3 and BD = 4x + 5 to get those answers \(\color{#0cbb34}{\text{End of Quote}}\)

Shaypop:

16.02 and 17.68

Shaypop:

Those the answers right

Angle:

16 and 17 2/3

Shaypop:

So what's the answer

Angle:

I said what I said

Shaypop:

What's Ad

Angle:

\(\color{#0cbb34}{\text{Originally Posted by}}\) Angle AD = AB \(\color{#0cbb34}{\text{End of Quote}}\)

Shaypop:

Ad has a long line

Shaypop:

You supposed to subtract

Angle:

if it's a ... LONG line no... I don't think you would subtract

Shaypop:

What's ce

Angle:

\(\color{#0cbb34}{\text{Originally Posted by}}\) Angle CE = BD divided by 2 AE = BD divided by 2 \(\color{#0cbb34}{\text{End of Quote}}\)

Shaypop:

16 divided by 2

Angle:

BD = 17 2/3

Shaypop:

It has to be a decimal

Angle:

then divide and convert to a decimal

Angle:

if you are in geometry, this should be something you should already know how to do

Shaypop:

I don't

Angle:

17 2/3 divided by 2

Shaypop:

Can you give me the answers

Shaypop:

8.3

Shaypop:

8.83

Angle:

yes

Shaypop:

So ac is 8.83

Angle:

no AC = BD

Shaypop:

Ac is 17 2/3

Angle:

yes

Shaypop:

Ab is 16

Angle:

yes

Shaypop:

Ad is

Shaypop:

16 divided by 2

Angle:

AD = AB

Shaypop:

Is ad is 16

Angle:

yes

Shaypop:

Ae

Angle:

\(\color{#0cbb34}{\text{Originally Posted by}}\) Angle CE = BD divided by 2 AE = BD divided by 2 \(\color{#0cbb34}{\text{End of Quote}}\)

Shaypop:

Is 8

Angle:

no

Shaypop:

Idk

Shaypop:

93 divided by 2

Angle:

no

Shaypop:

17.68

Angle:

divide that by 2

Shaypop:

8.84

Angle:

ok

Shaypop:

Bac

Angle:

\(\color{#0cbb34}{\text{Originally Posted by}}\) Angle angle BAC = 90 divided by 2 angle BDC = 90 divided by 2 \(\color{#0cbb34}{\text{End of Quote}}\)

Shaypop:

45

Angle:

ok